OpenStudy (anonymous):

Simplify by rationalizing the denominator (need the steps please.)Thanks y/(√7+√3)

OpenStudy (anonymous):

http://www.twiddla.com/563490

OpenStudy (anonymous):

multiply top and bottom by "conjugate" of bottom \[\frac{y}{\sqrt{7}+\sqrt{3}}\times \frac{\sqrt{7}-\sqrt{3}}{\sqrt{7}-\sqrt{3}}\]

OpenStudy (anonymous):

you get \[\frac{y(\sqrt{7}-\sqrt{3}}{7-3}=\frac{y(\sqrt{7}-\sqrt{3})}{4}\]

OpenStudy (anonymous):

you still have the radicals, but now they are in the numerator. this is why it is called "rationalizing the denominator". the second expression, however, is in no way "simpler" than the first.

OpenStudy (anonymous):

Did you get that levie

OpenStudy (anonymous):

yes